OpenStudy (anonymous):

what is the lcm for 6(y-1), 9-9y @tgawade @satellite73

OpenStudy (anonymous):

Do you know what is LCM ?

OpenStudy (anonymous):

least common multiple

OpenStudy (anonymous):

Correct !! So we have 6 * (y-1) & if we factor out -9 from 9 - 9y what is remain?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

please check it 9 - 9y so -9 (y - 1) right ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So we have 6 *(y-1) & -9 * (y - 1) so what would be the lcm?

OpenStudy (anonymous):

Here also find lcm of 6 & -9

OpenStudy (anonymous):

lcm = common factor * uncommon factor So can you tell me what are the factors of 6 = ---------------? & 9 = ---------------?

OpenStudy (anonymous):

6 :) 1,2,3 9:) 1,3,3 ritte

OpenStudy (anonymous):

So common factor = ----------------? & uncommon factor =--------------------?

OpenStudy (anonymous):

commom factor = 3 & uncommon = 3*2 So LCD = 3*3*2 = 18 okay ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

LCM=18

OpenStudy (anonymous):

So - 18*(y-1) is lcm

OpenStudy (anonymous):

we forgot -ve sign